OpenStudy (heart_offire):

Simplify. (3x^3 - 7x^2 + 12) - (3x^3 + 6x^2 + 10x) A) -x^2 + 10x + 12 B) -13x^2 - 10x + 12 C) -6x^3 + x^2 - 10x - 12 D) -9x^6 - 42x^4 - 10x + 12

OpenStudy (lolacole12):

oh Let's simplify step-by-step. 3x3−7x2+12−(3x3+6x2+10x) Distribute the Negative Sign: =3x3−7x2+12+−1(3x3+6x2+10x) =3x3+−7x2+12+−1(3x3)+−1(6x2)+−1(10x) =3x3+−7x2+12+−3x3+−6x2+−10x Combine Like Terms: =3x3+−7x2+12+−3x3+−6x2+−10x =(3x3+−3x3)+(−7x2+−6x2)+(−10x)+(12) =−13x2+−10x+12 Answer: =−13x2−10x+12

OpenStudy (heart_offire):

Why is there a negative 1 being distributed?